OpenStudy (anonymous):

Five cards are drawn from a deck of cards; Each card is replaced after being drawn. What is the probability of the following event: P(3 and 3 and 3 and 2 and 2)?

hartnn (hartnn):

can u just find the probability P(3) and P(2) ?

OpenStudy (anonymous):

I think that's what it's asking

hartnn (hartnn):

ok, how many cards are there in the deck ?

OpenStudy (anonymous):

I have no clue :/

hartnn (hartnn):

ohh.... there are total, 52 cards in a deck. 13 of clubs 13 of hearts 13 of spades 13 of diamonds

OpenStudy (anonymous):

Okay

hartnn (hartnn):

now u are drawing 5 cards from 52 cards. when u draw one card , the probability of card being 2 is 4/52 because there are 4 '2' in the deck. got this ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

similarly the probability of drawing a '3' is 4/52=1/13. so u have P(2)=1/13 and P(3) = 1/13 ok with this ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

since the events of drawing 2 and drawing 3 are independent. we have:P(3 and 3 and 3 and 2 and 2) = P(3)P(3)P(3)P(2)P(2) ok with this next step ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

so :P(3 and 3 and 3 and 2 and 2) = (1/13)*(1/13)*(1/13)*(1/13)*(1/13)=(1/13)^5 got this ?

OpenStudy (anonymous):

Okay, Yes.

hartnn (hartnn):

thats your final answer: (1/13)^5 or 1/(13)^5 or 1/(371293)

OpenStudy (anonymous):

Thank you! :) I don't get how you got 1/13 though?

hartnn (hartnn):

u got how i wrote 4/52 ?

OpenStudy (anonymous):

Yes, because there's 4 '2' in a deck. What does that mean though? 4 what in a deck?

hartnn (hartnn):

every deck contains 52 cards. 4 suits, which i mentioned. each suit has 13 cards from A,2,3,4,5,6,7,8,9,10,J,Q,K hence there will be 4 '2's of each suit.

hartnn (hartnn):

*there will be 4 '2's, one of each suit.

hartnn (hartnn):

hence the probability of selecting a '2' is 4/52 = 1/13

OpenStudy (anonymous):

Alright I got it now. Thank you :)

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

I might post another one that i'm a little condfused on, Can you help me?

hartnn (hartnn):

sure :)